LA= 1/2(36)(3√5)
can you please help and explain to me how to solve this? Thanks
Monika

Answer by rothauserc(4718)

We are given:
LA= 1/2(36)(3√5)
divide 36 by 2 and we get
LA=18(3sqrt(5))
multiply 18*3 and we get
LA=54sqrt(5)
sqrt(5) is 2.24, so we have
LA=54*2.24= 120.75